| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  1  
 | 
Wed Jun 15 13:11:38 2022 | 
Guy Leckenby | DAQ | Signal Diagram | Attached is the signal mapping that will be 
used for the BEAVERTAIL test.
 
 | 
    | 
| 
  2  
 | 
Fri Jun 17 15:32:47 2022 | 
Guy Leckenby | Electronics | Preamp Settings | Bias Select: In general, higher bias resistors 
-> lower electronic noise. However, higher 
resistance = larger voltage drop due to leakage | 
    | 
| 
  3  
 | 
Mon Jun 20 10:36:56 2022 | 
Guy Leckenby | DAQ | DAQ Startup Commands | Three terminals/processes are required to 
record data on a local machine.
 
1. log in x86l-131. 
 | 
   | 
| 
  4  
 | 
Tue Jun 21 02:37:07 2022 | 
Guy Leckenby, Iris Dillmann, Chris Griffin | General | Run notes | Beam: 208Pb(82+), 269.0 MeV/u, ~2*10^7 ions 
in ESR
 
 | 
      | 
| 
  5  
 | 
Wed Jun 22 18:01:18 2022 | 
Guy Leckenby | Detectors | Detector Arrangment | Attached is the detector arrangement for the 
3 setups used during the tests. | 
    | 
| 
  6  
 | 
Wed Jun 22 19:14:12 2022 | 
Guy Leckenby | DAQ | FBEX Script Choices | Here were the f_user.c choices we implemented 
for this experiment.
 
 | 
    | 
| 
  7  
 | 
Tue Jun 28 17:42:53 2022 | 
Guy Leckenby | DAQ | Cabling Changes | In the previous signal map, I incorrectly 
mapped the DSSD channels. To correctly read 
out the DSSD for the new test, we will swap | 
    | 
| 
  8  
 | 
Fri Jul 29 15:33:08 2022 | 
Guy Leckenby | Electronics | Results of Noise Testing | Uwe and I were able to get some noise testing 
done before the ESR was closed with asbestos 
exposure. We used the new flange so didn't | 
     | 
| 
  9  
 | 
Mon Nov  7 14:14:38 2022 | 
Chris Griffin, Guy Leckenby | Electronics | MSI-8 module details | MSI-8 module details.
 
 
MMPR1 card configurations:
 | 
   | 
| 
  10  
 | 
Fri Nov 11 10:27:59 2022 | 
Chris Griffin, Guy Leckenby | Detectors | Current leakage currents | Attached are photos of leakage currents for 
the current GAGG and BGO setups.
 
This setup uses the new flange (blue gas | 
     | 
| 
  11  
 | 
Tue Nov 15 14:30:25 2022 | 
Chris Griffin | Electronics | Old flange noise | Last week we tested both the new and old flanges 
with the GAGG housing set up described in 
the previous ELog entry.
 | 
    | 
| 
  12  
 | 
Thu Nov 17 13:42:16 2022 | 
Chris Griffin | Detectors | Detector config changes | I have changed the configuration of Housing 
2 (BGO) to account for the detector being 
fixed (bond wires) and provide a "plug-and-play" | 
   | 
| 
  14  
 | 
Mon May 29 12:23:48 2023 | 
Guy Leckenby, Chris Griffin | Electronics | Old flange noise re-tests | Whilst constructing a new flange to replace 
the old one, we did some tests again to compare 
the performance of new vs old flange.
 | 
        | 
| 
  15  
 | 
Thu Jun  1 09:39:24 2023 | 
Guy Leckenby, Chris Griffin | Electronics | New new flange test | We tested the new new flange using the same 
setup as in ELOG entry 14. The noise level 
seemed to be somewhere between the middle | 
      | 
| 
  16  
 | 
Thu Jun  1 10:42:31 2023 | 
Guy Leckenby | Electronics | Summary of Preamp Testing | MSI-8 and MPR-16 preamps have nominal differential 
output ranges of +/-1.8V and +/-1.2V respectively 
at maximum input (ie 5GeV).
 | 
    | 
| 
  17  
 | 
Tue Dec 12 12:57:16 2023 | 
Jan | DAQ | test lmd-files with pulser | We took a few lmd-files with different settings 
of the febex DAQ to provide data for go4 
development, when the DAQ is not running.
 | 
   | 
| 
  18  
 | 
Mon Mar 24 14:44:14 2025 | 
C Griffin, M Cockshutt, I Dillmann | General | April 2025 experiment setup | Further PLEIADES tests are to be carried out 
during E018 and E203. This will be an important 
test of PLEIADES operation in isochronous | 
   | 
| 
  19  
 | 
Fri Mar 28 13:31:53 2025 | 
CG, ID, MC, JG | General | Dual PLEIADES installation at ESR | PLEIADES detectors were installed in both 
the inside and outside pockets after the 
dipole behind the gas jet.
 | 
       | 
| 
  20  
 | 
Fri Mar 28 20:45:08 2025 | 
 CG | Electronics | MSI-8 configurations | The MSI-8 modules were reconfigured to properly 
instrument the current setup.
 
 | 
        | 
| 
  21  
 | 
Fri Mar 28 22:45:24 2025 | 
CG | Detectors | Bias problems | After testing the biasing scheme for all detectors, 
two problems emerge.
 
 | 
     | 
| 
  22  
 | 
Sat Mar 29 15:51:46 2025 | 
CG, JG, MC, ID | General | DAQ setup | ********************************************
 
** Setup of DAQ x86l-131 (outside pocket) 
**
 | 
   | 
| 
  23  
 | 
Sat Mar 29 17:19:46 2025 | 
CG, JG, MC, ID | General | Sample data file | Data files written to /mnt/raw.data/e0203_e0018.
 
 
From MBS computers:
 | 
   | 
| 
  24  
 | 
Sat Mar 29 17:53:13 2025 | 
CG, ID, MC | Electronics | Bias remote control | The MHV modules are set up for remote control. 
This can be accessed via the mesycontrol 
software running on x86l-128.
 | 
   | 
| 
  25  
 | 
Mon Mar 31 17:55:16 2025 | 
JG | DAQ | DAQ overview | The DAQ is running of 4 different nodes/PCs, 
that can be accessed by ssh using the user 
litv-exp.
 | 
   | 
| 
  26  
 | 
Mon Mar 31 18:21:37 2025 | 
JG | DAQ | ESR scaler mapping | This is the list of scaler channels used at 
the VME ESR DAQ on x86l-204:
 
 | 
   | 
| 
  27  
 | 
Mon Mar 31 18:35:46 2025 | 
JG | DAQ | lmd files on ESR NAS | The files of each MBS node or the timesorter 
are written to the ESR NAS by the RFIO server.
 
The server is running on the virtual machine | 
   | 
| 
  28  
 | 
Tue Apr  1 17:18:10 2025 | 
ID, MC, CJG | Detectors | Both PLEIADES detectors in beam | We started inserting the inside and outside 
detector, see
 
https://elog.gsi.de/esr/E0203_E0018/46
 | 
   | 
| 
  30  
 | 
Tue Apr  1 22:41:52 2025 | 
CG | Detectors | Climbing leakage currents | The experiment has been running stably during 
the afternoon.
 
DAQ rates have remained constant and the | 
    | 
| 
  31  
 | 
Tue Apr  1 23:00:41 2025 | 
CG | General | Run stopped and detectors moved back | Due to the continued rise in leakage currents, 
I have stopped the run, moved the detectors 
completely back and turned off the 
 | 
   | 
| 
  32  
 | 
Thu Apr  3 13:09:28 2025 | 
JG | DAQ | 194Pb MBS settings | The settings of the MBS daq during the first 
data taking with fragments around 194Pb (W.Korten 
experiment) was not entirely correct:
 | 
   | 
| 
  33  
 | 
Thu Apr  3 13:24:31 2025 | 
CG | Detectors | Bias problem diagnosing | During an access break to the ESR, I made 
a few tests with the bias scheme to see if 
we could better understand some of the weird | 
   | 
| 
  34  
 | 
Thu Apr  3 14:53:13 2025 | 
ID | Detectors | Link to leakage current Google doc | Please enter the leakage currents from the 
mesycontrol every ~1h into this Google document:
 
 | 
   | 
| 
  35  
 | 
Thu Apr  3 17:37:04 2025 | 
CJG, GL, MC | General | Detectors in position, taking data | The inner detector was moved into position.
 
The device control for the outer detector 
is not responding and I am not an expert/don't | 
   | 
| 
  36  
 | 
Fri Apr  4 16:12:36 2025 | 
CJG, GL, MC, JG, ID | General | Moving PLEIADES in | 186Hf now stored stably in the ring.
 
 
Will move both detectors in to position.
 | 
   | 
| 
  37  
 | 
Fri Apr  4 16:27:11 2025 | 
CJG, GL, MC, JG, ID | General | PLEIADES in position | Inner PLEIADES detector moved into position 
at +2mm.
 
Again, the outside pocket cannot be moved.
 | 
   | 
| 
  38  
 | 
Fri Apr  4 21:15:13 2025 | 
CJG, GL, MC, ID | General | Changes to setup | After talking to T. Davinson this afternoon 
a few things became clearer.
 
 | 
   | 
| 
  39  
 | 
Sat Apr  5 12:28:20 2025 | 
CG, GL, ID | General | Detector moved in | Inside pocket moved in to 2mm
 
 
run0024 started @ 12:27:50 | 
   | 
| 
  40  
 | 
Sat Apr  5 12:38:49 2025 | 
CG | General | Current status | Given that I hope we are now moving towards 
stable operation, here is the current status 
of detectors for run0024 on...
 | 
   | 
| 
  41  
 | 
Sat Apr  5 12:54:20 2025 | 
CG, MC | Detectors | Bias curves | Bias curves were completed for most detectors 
yesterday evening.
 
 | 
       | 
| 
  42  
 | 
Sat Apr  5 13:20:10 2025 | 
ID, CG, GL | General | MHV-4 max current | The leakage currents are 6-7 micro-amps for 
the Si pads at the moment. The max current 
for the MHV-4 is 20 microamps PER CHANNEL, | 
   | 
| 
  43  
 | 
Sat Apr  5 14:11:55 2025 | 
CG, GL, ID | General | Run stopped - detectors toast? | run0034 stopped at 14:11
 
 
Leakage currents are increasing and once | 
   | 
| 
  44  
 | 
Sat Apr  5 15:19:11 2025 | 
Team PLEIADES | General | Run started | 15:14 Started run #35: inside pocket at 20 
mm.
 
Trigger mask changed. Gated only in ch8 in | 
   | 
| 
  45  
 | 
Sat Apr  5 17:49:06 2025 | 
CG, GL, ID | Detectors | Outside detector moved in  | After some Sergey-magic the outside pocket 
moves again. We slowly moved the outside 
detector in, we are clipping the beam at | 
   | 
| 
  46  
 | 
Mon Apr  7 20:00:13 2025 | 
Guy | General | Results of scope test | At 10:50am today, we connected a remote Picoscope 
to the first 4 n-sides, one end of both resistive 
chains, and both crystal channels to directly | 
     | 
| 
  47  
 | 
Tue Apr  8 09:53:45 2025 | 
GL, ID, CG, MC |   | Moving in Outside Detector | At 9:53 we moved the outside detector in to 
take a short data run. We went -110 mm compared 
to a fully retracted position -126 mm. Data | 
   | 
| 
  48  
 | 
Tue Apr  8 11:37:30 2025 | 
GL, ID, CG, MC, JG |   | Moving in Outside Detector | We start recording data in run 39 and the 
outside detector is moved to -100 mm at 11:36. 
At 11:59 the detector was moved back to the | 
   | 
| 
  49  
 | 
Wed Apr  9 10:38:05 2025 | 
CG | General | End of experiment | Final injection stored in ring around 10:00 
and will be stored for ~1h.
 
Both detectors are in the parked position | 
   | 
| 
  50  
 | 
Wed Apr  9 11:02:35 2025 | 
Team PLEIADES | Detectors | Beam stopped, voltages off | 11:01 All voltages turned off. Remote control 
disabled. | 
   | 
| 
  51  
 | 
Thu Apr 10 14:58:02 2025 | 
CG, MC | General | Post-run bias curves | BGO detector assembly originally on flange 
A.
 
GAGG  "                        "    flange | 
   | 
| 
  52  
 | 
Fri Apr 11 14:07:39 2025 | 
CG | General | End of run summary | All detector components are now in safe storage 
in Jan's container.
 
 | 
    |